OpenStudy (anonymous):

PLEASE HELP! Prove that if \[ b_n\] converges to \[ B\] and \[B \neq 0\], then there is a positive real number \[ M\] and a positive integer \[ N \] such that if \[n \geq N\], then \[\left | b_n \right |\geq M\]

OpenStudy (anonymous):

Wait.. it converges to \(nB\)? Does \(n\) head toward infinity when we take the limit?

OpenStudy (anonymous):

I got it fixed. It was a typo

OpenStudy (zzr0ck3r):

Suppose \(b_n \rightarrow b>0\) there exists a \(n_0\in N\) s.t. \(\frac{1}{n_0}<b\) Let \(\epsilon =\frac{1}{n_0}\) \(\exists N \ \forall \ n\ge N \ |b_n-b|<\frac{1}{n_0}\\\implies -\frac{1}{n_0}+b<b_n\) note that \(b-\frac{1}{n_0}>0\). So now we need a case for \(b<0\)

OpenStudy (zzr0ck3r):

Do you follow this?

OpenStudy (anonymous):

Yes am following

OpenStudy (zzr0ck3r):

The negative case should be similar, you can bound it away from 0, on the negative side, then look at the negative terms

OpenStudy (anonymous):

Is there a specific theorem tied to this?

OpenStudy (zzr0ck3r):

nope, its useful to prove many theorems.